4.(15 pts) For the following data used to predict number of yearly doctors visits from number of pets owned: Pets Doc Visits 6 11 12 2 7 12 a. Find the slope b. Find the intercept c. Write out the equation of the regression line and predict the number of yearly doctors visits for a person who owns 0 pets. 5. (12 pts) Using the data from question 4, assume you are now interested in predicting number of pets owned from number of yearly doctors visits: a. Find the slope b. Find the intercept d. Write out the equation of the regression line and predict the number of pets owned for a person who visits the doctor 50 times a year.
Answer #1

4. -0.75 5.25 -4.75 0.25 0.5625 27.5625 22.5625 0.0625 -3.375 -23.625 26.125 1.375 -5.5 5.5 SS 50.75 SP: 0.5

Sum of X = 27
Sum of Y = 26
Mean X = 6.75
Mean Y = 6.5
Sum of squares (SSX) = 50.75
Sum of products (SP) = 0.5

Regression Equation = ŷ = bX + a

Slope=b = SP/SSX = 0.5/50.75 = 0.00985

Intercept=a = MY - bMX = 6.5 - (0.01*6.75) = 6.4335

ŷ = 0.00985X + 6.4335